# TopoBenchmarkX: A Framework for Benchmarking Topological Deep Learning

### v1

@@ Appendix C Training details and additional results
In this section, we delve into the details of our hyperparameter search methodology, optimization strategy, computational resources utilized for conducting the experiments, and present additional results and analyses.
C.1 Experiments’ configuration and model execution
@@
To automate the configuration of the TopoBenchmarkX modules we utilize hydra package [Yadan, 2019]. In particular, we rely on hierarchical configuration groups and registers to facilitate the easy use of the library. Specifically, there is no need to meticulously select each module to execute the pipeline in any of the domains. Simply choosing a dataset and a model automatically configures a full default pipeline, eliminating the need for manual intervention.
TopoBenchmarkX also automates the model execution and learning through the lightning library [Lisa and Bot, 2017], orchestrating the training, validation and testing processes –as well as handling any logging and callbacks.
C.2 Hyperaparameter search
@@
Five splits are generated for each dataset to ensure a fair evaluation of the models across domains. Each split comprises 50% training data, 25% validation data, and 25% test data. An exception is made for the ZINC dataset, where pre-defined splits are used [Irwin et al., 2012].
@@
Each model in every domain has numerous specific hyperparameters that can be tuned to enhance performance. TNNs, in particular, have multiple additional parameters that can potentially improve results. To avoid the combinatorial explosion of possible hyperparameter sets, the search space is limited to the hyperparameters common across all models. The grid-search strategy is employed to find the optimal hyperparameters for each model and dataset. Specifically, the encoder hidden dimension is varied over [32, 64, 128], the encoder dropout is varied over [0.25, 0.5], the number of backbone layers is varied over [1, 2, 3, 4], the learning rate is varied over [0.01, 0.001], and the batch size is varied over [128, 256]. Additionally, for the cellular and simplicial domains, the readout type is varied over [direct readout (DR), signal down-propagation (SDP)]. If the model cannot fit into memory, the batch size, encoder hidden dimension, and number of backbone layers are decreased until the model training fits into the GPU memory.
@@
For node-level task datasets, validation is conducted after each training epoch, continuing until either the maximum number of epochs is reached or the optimization metric fails to improve for 50 consecutive validation epochs. The minimum number of epochs is set to 50. Conversely, for graph-level tasks, validation is performed every 5 training epochs, with training halting if the performance metric does not improve on the validation set for the last 10 validation epochs. To optimize the models, torch.optim.Adam is combined with torch.optim.lr_scheduler.StepLR wherein the step size was set to 50 and the gamma value to 0.5. In total more than 100,000 runs have been performed to obtain the results. The optimal hyperparameter set is generally selected based on the best average performance over five validation splits. For the ZINC dataset, five different initialization seeds are used to obtain the average performance.
@@
The hyperparameter search is executed on a Linux machine with 256 cores, 1TB of system memory, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory.
C.3 Results
@@
Table 3 additionally presents the results for the CCXN and SCCN networks, which on average, perform slightly worse than the other models. As shown in Table 4, the CCXN network performs better when using the SDP readout strategy, although the performance improvements are not as significant as those achieved by the CWN network with the same strategy. The SCCN model benefits more from utilizing the SDP readout compared to other simplicial domain models (SCN and SCCNN), showing performance improvements in 9 out of 21 cases, while SCCNN and SCN show improvements in only 3 and 5 cases, respectively. Overall, cellular models demonstrate performance enhancements on 15, 19, and 8 datasets for CCXN, CWN, and CCCN, respectively, with the SDP readout strategy. Conversely, simplicial domain models achieve 9, 3, and 5 performance improvements for SCCN, SCCNN, and SCN, respectively, with the same SDP strategy. Based on this analysis and the findings in Section 5.3, it is recommended for the new people in the TDL to opt for the SDP readout strategy if they prefer not to delve into the intricate details of interior propagation in cellular and simplicial models.
@@
Note that for demonstration purposes, only one fixed lifting is considered for lifting graphs to each of the considered topological domains, leaving out of the scope of this paper the analysis of the optimal lifting in each particular scenario. [^6] Note that learnable liftings can further optimize the predictive capacity of higher-order networks. In particular, the clique complex lifting is applied to lift the graph to a simplicial domain. Cycle-based lifting is employed to lift the graph into the cellular domain, and $k$-hop lifting is utilized to lift the graph into a hypergraph, where $k=1$. Projection lifting is applied to lift the features, specifically, the features of the $n-1$-cell projected to the $n$-cell by multiplying $n-1$-cell representation to corresponding incidence matrices.
@@
Acronym Neural network name Reference Graph neural networks GAT Graph attention network Veličković et al. [2018] GIN Graph isomorphism network Xu et al. [2019] GCN Semi-Supervised Classification with Graph Convolutional Networks Kipf and Welling [2016] Simplicial complexes SAN Simplicial Attention Neural Networks Giusti et al. [2022] SCCN Efficient Representation Learning for Higher-Order Data with Simplicial Complexes Yang et al. [2022] SCCNN Convolutional Learning on Simplicial Complexes Yang and Isufi [2023] SCN Simplicial Complex Neural Networks Ebli et al. [2020] Cellular complexes CAN Cell Attention Network Giusti et al. [2023b] CCCN Generalized simplicial attention neural networks  [^7] We report the results for the cellular domain of this implementation. Battiloro et al. [2023b] CXN Cell Complex Neural Networks Hajij et al. [2020] CWN Weisfeiler and Lehman Go Cellular: CW Networks Bodnar et al. [2021a] Hypergraphs AllSetTransformer You are AllSet: A Multiset Function Framework for Hypergraph Neural Networks Chien et al. [2021] EDGNN Equivariant Hypergraph Diffusion Neural Operators Wang et al. [2022] UniGNN UniGNN: a Unified Framework for Graph and Hypergraph Neural Networks Huang and Yang [2021]
Appendix D Higher-order domains
A higher-order domain is a generalization of a graph that captures both pairwise and higher-order relationships between entities Hajij et al. [2023]. In the study of topological domains, understanding the distinct properties of various higher-order domains is essential. Two key properties are set-type relations and hierarchical structures, represented by rank functions. Higher-order domains are also called topological domains.
A clarification on terminology. From a mathematical point of view, a higher-order domain is not necessarily a topological space. However, a higher-order domain consists of higher-order cells that capture information about the topology of the underlying domain. From this high-level perspective, higher-order domains (including hypergraphs) are cell structures that encode topological information. In such a context, the terms ‘higher-order domain’ and ‘topological domain’ are used interchangeably. Similarly, the terms ‘higher-order neural network’ and ‘topological neural network’ are used interchangeably.
Definition 4 (Set-type relation)
A relation in a higher-order domain is called a set-type relation if its existence is not implied by another relation in the domain.
Set-type relations emphasize the independence of connections within a domain, allowing for a flexible representation of complex interactions.
Definition 5 (Rank function)
A rank function on a higher-order domain $\mathcal{X}$ is an order-preserving function $rk\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$; i.e., $x\subseteq y$ implies $rk(x)\leq rk(y)$ for all $x,y\in\mathcal{X}$.
Rank functions introduce a hierarchical organization, facilitating the analysis of nested relationships and enabling pooling operations in deep learning by aggregating information across different levels of the hierarchy.
Over the past few years, four topological domains have found significant applications in TDL: hypergraphs, simplicial complexes, cell complexes, and combinatorial complexes. Note that currently, TopoBenchmarkX library supports simplicial complexes, cell complexes, and hypergraphs, while the rest will be added in future work.
D.1 Hypergraphs
Definition 6 (Hypergraph)
A hypergraph $\mathcal{H}$ on a nonempty set $\mathcal{V}$ is a pair $(\mathcal{V},\mathcal{E}_{\mathcal{H}})$, where $\mathcal{E}_{\mathcal{H}}$ is a non-empty subset of the powerset $\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$ of the nodes. Elements of $\mathcal{E}_{\mathcal{H}}$ are called hyperedges.
Hypergraphs generalize graphs by allowing edges (called hyperedges) to connect any number of nodes, representing set-type relations without hierarchical structure.
D.2 Abstract simplicial complexes
Definition 7 (Abstract simplicial complex)
An abstract simplicial complex (SC) in a non-empty set $S$ is a pair $SC=(S,\mathcal{X})$, where $\mathcal{X}$ is a subset of $\mathcal{P}(S)\setminus\{\emptyset\}$ such that $x\in SC$ and $y\subseteq x$ imply $y\in SC$. Elements of $\mathcal{X}$ are called simplices.
Simplicial complexes represent hierarchical structures where each simplex (relation) includes all its subsets, but they do not represent set-type relations independently.
D.3 Regular cell complexes
Definition 8 (Regular cell complex)
A regular cell complex is a topological space $S$ with a partition into subspaces (cells) $\{x_{\alpha}\}_{\alpha\in P_{S}}$, where $P_{S}$ is an index set, satisfying the following conditions:
  1. 1.
  2. 2.
    For each $\alpha\in P_{S}$, there exists a homeomorphism $\psi_{\alpha}$, called an attaching map, from $x_{\alpha}$ to $\mathbb{R}^{n_{\alpha}}$ for some $n_{\alpha}\in\mathbb{N}$, called the dimension $n_{\alpha}$ of cell $x_{\alpha}$.
  3. 3.
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a union of finitely many cells, each having dimension less than the dimension of $x_{\alpha}$.
Cell complexes provide a hierarchical interior-to-boundary structure with clear topological and geometric interpretations, but they are not based on set-type relations.
D.4 Combinatorial complexes
Definition 9 (Combinatorial complex)
A combinatorial complex (CC) is a triple $(\mathcal{V},\mathcal{X},\operatorname{rk})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{X}$ of $\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$, and a function $\operatorname{rk}\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$ with the following properties:
  1. 1.
    For all $v\in\mathcal{V}$, $\{v\}\in\mathcal{X}$ and $\operatorname{rk}(v)=0$,
  2. 2.
    The function $\operatorname{rk}$ is order-preserving, which means that if $x,y\in\mathcal{X}$ satisfy $x\subseteq y$, then $\operatorname{rk}(x)\leq\operatorname{rk}(y)$.
Combinatorial complexes combine hierarchical structure with set-type relations, allowing for a flexible and comprehensive representation of higher-order networks.
Appendix E Topological neural networks
Roughly speaking, a topological neural network (TNN) is a neural network that process data on a topological domain. In this section we review the basics of TNNs using the higher-order message passing paradiam introduced in Hajij et al. [2023]. All networks we experiment on in our package fall as a special case of the definition we provide here.
First, we introduce the concept of $k$-cochain spaces, which are essential building blocks for topological neural networks.
Definition 10 (k-cochain spaces)
Let $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ be the $\mathbb{R}$-vector space of functions $\mathbf{H}_{k}\colon\mathcal{X}^{k}\to\mathbb{R}^{d}$ for a rank $k\in\mathbb{Z}_{\geq 0}$ and dimension $d$. $d$ is called the data dimension. $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ is called the $k$-cochain space. Elements $\mathbf{H}_{k}$ in $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ are called $k$-cochains or $k$-signals.
Cochain represent the input topological data (feature vectors) defined on the topological domain. To describe how data is exchanged within a topological domain, we recall the concept of neighborhood functions.
Definition 11 (Neighborhood function)
Let $S$ be a nonempty set. A neighborhood function on $S$ is a function $\mathcal{N}\colon S\to\mathcal{P}(\mathcal{P}(S))$ that assigns to each point $x$ in $S$ a nonempty collection $\mathcal{N}(x)$ of subsets of $S$. The elements of $\mathcal{N}(x)$ are called neighborhoods of $x$ with respect to $\mathcal{N}$.
Definition 12 (TNN)
Let $\mathcal{X}$ be a topological domain. Let $\mathcal{C}^{i_{1}}\times\mathcal{C}^{i_{2}}\times\ldots\times\mathcal{C}^{i_{m}}$ and $\mathcal{C}^{j_{1}}\times\mathcal{C}^{j_{2}}\times\ldots\times\mathcal{C}^{j_{n}}$ be a Cartesian product of $m$ and $n$ cochain spaces defined on $\mathcal{X}$. A topological neural network (TNN) is a function of the form
$TNN\colon\mathcal{C}^{i_{1}}\times\mathcal{C}^{i_{2}}\times\ldots\times\mathcal{C}^{i_{m}}\longrightarrow\mathcal{C}^{j_{1}}\times\mathcal{C}^{j_{2}}\times\ldots\times\mathcal{C}^{j_{n}}.$
A TNN takes as input a vector of cochains $(\mathbf{H}_{i_{1}},\ldots,\mathbf{H}_{i_{m}})$ and produces as output a vector of cochains $(\mathbf{K}_{j_{1}},\ldots,\mathbf{K}_{j_{n}})$. This framework allows for the processing of complex topological data through neural network operations.
Finally, we introduce the concept of higher-order message passing, which is the mechanism by which TNNs can exchange information across different cells in the topological domain:
E.1 Higher-order message passing
Higher-order message passing refers to a computational framework that involves exchanging messages among entities and cells in a higher-order domain using a set of neighborhood functions.
Definition 13 (Higher-order message passing on a topological domain)
Let $\mathcal{X}$ be a topological domain, and let $\mathcal{N}=\{\mathcal{N}_{1},\ldots,\mathcal{N}_{n}\}$ be a set of neighborhood functions defined on $\mathcal{X}$. Consider a cell $x$ and another cell $y\in\mathcal{N}_{k}(x)$ for some $\mathcal{N}_{k}\in\mathcal{N}$. A message $m_{x,y}$ between cells $x$ and $y$ is a computation depending on these two cells or on the data they support. Let $\mathcal{N}(x)$ denote the multi-set $\{\!\!\{\mathcal{N}_{1}(x),\ldots,\mathcal{N}_{n}(x)\}\!\!\}$, and let $\mathbf{h}_{x}^{(l)}$ represent the data supported on the cell $x$ at layer $l$. Higher-order message passing on $\mathcal{X}$, induced by $\mathcal{N}$, is defined through the following update rules:
$m_{x,y}$ $=\alpha_{\mathcal{N}_{k}}(\mathbf{h}_{x}^{(l)},\mathbf{h}_{y}^{(l)}),$ (1)
$m_{x}$ $=\bigotimes_{\mathcal{N}_{k}\in\mathcal{N}}m_{x}^{k},$ (3)
$\mathbf{h}_{x}^{(l+1)}$ $=\beta(\mathbf{h}_{x}^{(l)},m_{x}).$ (4)
Here, $\bigoplus$ is a permutation-invariant aggregation function called the intra-neighborhood aggregation of $x$, and $\bigotimes$ is an aggregation function called the inter-neighborhood aggregation of $x$. The functions $\alpha_{\mathcal{N}_{k}}$ and $\beta$ are differentiable functions.
To summarize:
  • $m_{x,y}$ is the message computed from $x$ to $y$ using the function $\alpha_{\mathcal{N}_{k}}$.
  • $m_{x}^{k}$ aggregates all messages from the neighbors $y$ in the neighborhood $\mathcal{N}_{k}(x)$ using the intra-neighborhood function $\bigoplus$.
  • $m_{x}$ further aggregates these results across all neighborhoods $\mathcal{N}_{k}\in\mathcal{N}$ using the inter-neighborhood function $\bigotimes$.
  • $\mathbf{h}_{x}^{(l+1)}$ updates the data on cell $x$ by combining its current data $\mathbf{h}_{x}^{(l)}$ with the aggregated message $m_{x}$ using the function $\beta$.
E.2 Information about neural network models
In this section, we review the architectures of topological neural networks (TNNs) applied to hypergraphs, simplicial complexes, and cellular complexes; for a more comprehensive survey on TNN architectures, refer to Papillon et al. [2023b].
E.2.1 Hypergraphs neural networks
Hypergraph neural networks (HGNNs) have been extensively researched and surveyed. Early methods often reduced hypergraphs to graphs using algorithms such as clique-expansion, which compromised performance by losing structural information Dong et al. [2020]. Current models preserve hypergraph structures and typically employ a two-phase message-passing scheme, transferring information from nodes to hyperedges and then back to nodes.
Standard message passing. Models such as HyperSage, HGNN, Dynamic Hypergraph Neural Network, and UniGCN use standard message passing Wang et al. [2019], Liu et al. [2021], Huang and Yang [2021], Chien et al. [2021]. Notably, some models like UniGCN incorporate fixed weight matrices on top of learnable weights. Other approaches, such as AllSet and EHNN, enhance expressiveness using multi-set functions and sparse symmetric tensors, respectively Chien et al. [2021], Wang et al. [2022].
Attention-based message passing. Models like Dynamic Hypergraph Convolutional Network (SAN), Less is More: Hypergraph Attention Network (SAT), and UniGAT integrate attention mechanisms into the two-phase scheme. Transformer-based variants and multi-head attention are also employed to further enhance performance Wang et al. [2019], Liu et al. [2021].
E.2.2 Simplicial complexes neural networks
Initial research on simplicial complexes focused on signal processing. HodgeNet was one of the first models to learn convolutions on edge features, while subsequent models like SNN and SCCONV generalized convolutions to higher-order features Bodnar et al. [2021a], Bunch et al. [2020], Yang and Isufi [2023].
Standard message passing. Simplicial neural networks (SNNs), SCCONV, and more recent models utilize various neighborhood schemes to enhance message passing flexibility Giusti et al. [2022], Ebli et al. [2020]. High Skip Networks (HSNs) introduce higher-order skip-connections to propagate features through multiple ranks Hajij et al. [2022a].
Attention-based message passing. Models such as SAN, SAT, and SGAT leverage attention mechanisms on simplicial domains, each utilizing unique neighborhood structures and attention coefficients Giusti et al. [2022], Hajij et al. [2022b], Battiloro et al. [2023b].
E.2.3 Cellular complexes neural networks
Building on signal processing advancements, Cell Complex Neural Networks (CXNs) Hajij et al. [2020] were introduced theoretically and later implemented in (CWN) Bodnar et al. [2021a]. The only attention-based cellular network that is Cell Attention Networks (CAN) introduced in Giusti et al. [2023b]. Recently, a cellular transformer architecture (CT) was introduced in Ballester et al. [2024].
Appendix F Descriptive summaries of datasets
Table 7 provides the statistics for each dataset lifted to three topological domains: simplicial complex, cellular complex, and hypergraph. The 00-cell, $1$-cell, $2$-cell, and $3$-cell represents the count of $n$-cell each dataset have after the topology lifting procedure. The clique complex lifting is applied to lift the graph to a simplicial domain with a maximum $n$ equal to three. Cycle-based lifting is employed to lift the graph into the cellular domain (maximum $n$ is equal to two), and $k$-hop lifting is utilized to lift the graph into a hypergraph, where $k=1$.

### v2

@@ Appendix C Further Experimental Details
This appendix provides details on the hyperparameter search methodology, optimization strategy, computational resources used for the experiments, and additional results and analyses.
C.1 Experiment Configuration and Model Execution
@@
To automate the configuration of TopoBench modules, the hydra package (Yadan, 2019) is employed. In particular, hierarchical configuration groups and registers facilitate easy use of the library: there is no need to meticulously select each module for any given domain. Simply choosing a dataset and a model automatically configures a full default pipeline, eliminating the need for manual intervention. Model execution and training are further automated by the lightning library (Lisa and Bot, 2017), which orchestrates training, validation, and testing while also handling logging and callbacks.
C.2 Hyperparameter Search
@@
Five splits are generated for each dataset to ensure a fair evaluation across domains, allocating 50% of data for training, 25% for validation, and 25% for testing. An exception is made for the ZINC dataset, which uses predefined splits (Irwin et al., 2012).
@@
Each model (in each domain) has numerous specific hyperparameters that can be tuned to enhance performance. TNNs, in particular, come with additional parameters that could further boost results. To avoid the combinatorial explosion of all possible hyperparameter sets, the search space is restricted to hyperparameters common across every model. A grid-search strategy is used to identify the optimal parameters for each model-dataset combination. Specifically, the encoder hidden dimension is varied over {32, 64, 128}, the encoder dropout over {0.25, 0.5}, the number of backbone layers over {1, 2, 3, 4}, the learning rate over {0.01, 0.001}, and the batch size over {128, 256}. For models in the cellular and simplicial domains, the readout type is also varied between direct readout (DR) and signal down-propagation (SDP). If a model exceeds available GPU memory, the batch size, encoder hidden dimension, and number of backbone layers are reduced until training becomes feasible.
@@
For node-level tasks, validation is conducted after each training epoch, continuing until either the maximum number of epochs is reached or the optimization metric fails to improve for 50 consecutive validation epochs; a minimum of 50 epochs is always enforced. For graph-level tasks, validation is performed every 5 epochs, halting early if validation performance fails to improve for 10 consecutive checks. The optimization uses torch.optim.Adam combined with torch.optim.lr_scheduler.StepLR, where the step size is 50 and $\gamma=0.5$. Over 100,000 runs have been executed to obtain the final results. In general, the best hyperparameter set is selected based on the highest average performance across the five validation splits. For ZINC, five different initialization seeds are used to obtain an average performance.
@@
All experiments are conducted on a Linux machine equipped with 256 CPU cores, 1TB of system memory, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory.
C.3 Descriptive Summaries of Datasets
Table 4 provides descriptive statistics for each dataset used in the numerical experiments (see Section 5 for details) after lifting to three topological domains: simplicial complexes, cellular complexes, and hypergraphs. The columns labeled 00-cell, $1$-cell, $2$-cell, and $3$-cell show the count of each $n$-cell in the resulting dataset. Specifically, a clique complex lifting is applied to obtain a simplicial domain with a maximum dimension of three, cycle-based lifting is used to obtain a cellular domain with a maximum dimension of two, and $k$-hop lifting (with $k=1$) is used to lift each graph into a hypergraph.
C.4 Additional Results
@@
Table 5 additionally presents results for the CCXN and SCCN networks, which on average perform slightly worse than other models. As shown in Table 6, the CCXN network performs better when using SDP readout, though not as dramatically as CWN under the same strategy. The SCCN model benefits more from SDP readout than other simplicial-domain models (SCN and SCCNN), showing improvements in 9 out of 21 cases, whereas SCCNN and SCN show improvements in 3 and 5 cases, respectively. Overall, cellular models demonstrate improved performance on 15, 19, and 8 datasets for CCXN, CWN, and CCCN, respectively, when using SDP. In contrast, simplicial models achieve 9, 3, and 5 improvements for SCCN, SCCNN, and SCN, respectively, with the same readout.
@@
Note that for demonstration purposes, only one fixed lifting is applied to transform graphs into each of the considered topological domains, leaving a domain-specific optimal lifting strategy beyond the scope of this paper. [^6] Learnable liftings may further optimize the predictive capacity of higher-order networks. Specifically, a clique complex is used for simplicial lifting, cycle-based lifting is used for cellular domains, and $k$-hop lifting (with $k=1$) is used for hypergraphs. Feature projection is also applied, where the $(n-1)$-cell features are multiplied by the corresponding incidence matrices to generate $n$-cell features.
@@
Acronym Neural network name Reference Graph neural networks GAT Graph attention network Veličković et al. (2018) GIN Graph isomorphism network Xu et al. (2019) GCN Semi-Supervised Classification with Graph Convolutional Networks Kipf and Welling (2016) Simplicial complexes SAN Simplicial Attention Neural Networks Giusti et al. (2022) SCCN Efficient Representation Learning for Higher-Order Data with Simplicial Complexes Yang et al. (2022) SCCNN Convolutional Learning on Simplicial Complexes Yang and Isufi (2023) SCN Simplicial Complex Neural Networks Ebli et al. (2020) Cellular complexes CAN Cell Attention Network Giusti et al. (2023) CCCN Generalized simplicial attention neural networks  [^7] We report the results for the cellular domain of this implementation. Battiloro et al. (2024b) CXN Cell Complex Neural Networks Hajij et al. (2020) CWN Weisfeiler and Lehman Go Cellular: CW Networks Bodnar et al. (2021a) Hypergraphs AllSetTransformer You are AllSet: A Multiset Function Framework for Hypergraph Neural Networks Chien et al. (2021) EDGNN Equivariant Hypergraph Diffusion Neural Operators Wang et al. (2022) UniGNN UniGNN: a Unified Framework for Graph and Hypergraph Neural Networks Huang and Yang (2021)